OpenStudy (anonymous):

Select the ordered pair from the choices below that is a solution to the following system of equations: 3x = y + 18 y + 7 = -30x a. (7, 3) b. (-2, 12) c. (1/3, -17) d. (-4, 2)

OpenStudy (callisto):

3x = y + 18 -(1) y + 7 = -30x => y + 7 = -10 (3x) -(2) Sub. (1) into (2) y + 7 = -10 ( y + 18 ) y + 7 = -10y - 180 11y = -187 y = ...?

OpenStudy (anonymous):

-17 ? @Callisto so my answer is C ?

OpenStudy (callisto):

Yes...